Consider a consumer whose utility function is given as: (x, y) = xy, where and denote the quantities of goods and consumed. The budget constraint faced by the consumer is: 4+ 8= 120, where 4 is the price of good x, 8 is the price of good and 120 is the income of the consumer.



a) Find the optimal quantities for and consumed by the consumer. Show your solution diagrammatically


b)Following on the answer in a, now assume that the price of good increases to 8. Find the new quantities consumed by the consumer


1
Expert's answer
2022-01-09T18:06:08-0500


U(X,Y)=XY

"P_x=4"

"P_y=8"

4x+8y=120

a) For Utility Maximization,

"\frac{Mu_x} {Mu_y} =\frac{P_x}{P_y}"

"\frac{4} {8} =\frac{P_x}{P_y}"

4x=8y

x=2y

y=0.5x

From the budget line we have;

4(2y)+8y=120

8y+8y=120

16y=120

y=7.5


So, consumption bundle =(15,7.5)



b) Increase in price of x to 8

"P_x=8"

"P_y=8"

8x+8y=120

a) For Utility Maximization,

"\frac{Mu_x} {Mu_y} =\frac{P_x}{P_y}"

"\frac{8} {8} =\frac{P_x}{P_y}"

"\frac YX =1"

From the budget line we have;

8(y)+8y=120

8y+8y=120

16y=120

y=7.5

8x+ 8(x)=120

8x+8x=120

16x=120

x= 7.5

So, consumption bundle =(7.5,7.5)